### v2

# Large Language Model Bias Mitigation from the Perspective of Knowledge Editing

Abstract
Existing debiasing methods inevitably make unreasonable or undesired predictions as they are designated and evaluated to achieve parity across different social groups but leave aside individual facts, resulting in modified existing knowledge. In this paper, we first establish a new bias mitigation benchmark BiasKE leveraging existing and additional constructed datasets, which systematically assesses debiasing performance by complementary metrics on fairness, specificity, and generalization. Meanwhile, we propose a novel debiasing method, Fairness Stamp (FAST), which enables editable fairness through fine-grained calibration on individual biased knowledge. Comprehensive experiments demonstrate that FAST surpasses state-of-the-art baselines with remarkable debiasing performance while not hampering overall model capability for knowledge preservation, highlighting the prospect of fine-grained debiasing strategies for editable fairness in LLMs.
1 Introduction
Pre-trained Large Language Models (LLMs) have demonstrated exceptional performance on many tasks (Devlin et al., 2018; Floridi & Chiriatti, 2020; Brown et al., 2020). However, the encoded social stereotypes and human-like biases inevitably cause undesired behaviors when deploying LLMs in practice (Zhao et al., 2019; Navigli et al., 2023; Sheng et al., 2021). Existing approaches to mitigate biases in LLMs are mainly categorized into: (1) Fine-tuning (Zmigrod et al., 2019; Webster et al., 2020; He et al., 2022; Liang et al., 2020; Lauscher et al., 2021), which includes techniques such as re-balanced corpus pre-training, contrastive learning, projection methods, and efficient parameter tuning. (2) Prompt-tuning (Guo et al., 2022; Yang et al., 2023; Li et al., 2023b; Dong et al., 2023), which involves creating prompts to address social biases.
Figure 1: (a) Expression towards different groups (e.g., mom/dad) does not necessarily constitute a bias. (b) Existing debiasing approaches usually equalize different groups, resulting in unreasonable predictions. (c) Our proposed method performs fine-grained calibration with biased knowledge, while maintaining the others.
However, existing techniques treat social groups as interchangeable (Gallegos et al., 2023) and neutralize protected attributes of different social groups in model inputs or outputs, while ignoring or concealing distinct mechanisms of different social groups (Hanna et al., 2020), as shown in Figure 1. Furthermore, existing debiasing evaluation metrics mainly focus on the degree of bias, but fail to measure whether the model retains its origin knowledge (Gallegos et al., 2023) of discerning reasonable disparities among different social groups.
To address these issues, we first establish a more comprehensive debiasing benchmark BiasKE by extending existing datasets with additional constructed data and evaluation metrics on fairness, specificity, and generalization. Moreover, we propose a novel method Fairness-Stamp (FAST) for editable bias mitigation. Instead of mitigating group biases indiscriminately, FAST operates fine-grained calibrations on individual biases, i.e., specific stereotyped statements toward a social group. Specifically, we first design a causal-tracing-based method to locate the decisive layer in LLMs responsible for biased predictions. Then we propose to add a lightweight modular network, which enables fine-grained and efficient debiasing of one or multiple individual biased knowledge, with objectives of bias mitigation and knowledge maintenance.
We evaluate FAST with comprehensive experiments on StereoSet (Nadeem et al., 2020b) and Crows-Pairs (Nangia et al., 2020), which are further extended as BiasKE for systematic evaluation. Results show that FAST achieves remarkable debiasing performance without compromising model capability. We extend FAST to larger models such as GPT-Neo and Llama to demonstrate the scalability in real-world applications. Additional experiments showcase the effectiveness on downstream tasks, continual bias mitigation, and lightweight optimization, with results and analysis in Appendix D.
2 BiasKE Benchmark Construction
Figure 2: An illustration of the construction of BiasKE.
In this section, we describe the procedures for establishing BiasKE, with an illustration in Figure 2. To better express a bias, we formalize the stereotype bias (e.g., Man is good at man) as a triplet $k=(s,r,o)$, where $s$ is the subject (i.e., Man), $o$ is the object (i.e., math), and $r$ is the relation between them (i.e., is good at), as inspired by Petroni et al. (2019). We collect social biases related to three domains (gender, race, and religion) from six existing datasets, as detailed in Appendix A.2.
Step1. Based on these social biases, we extract biased knowledge pairs ${(k_{1},k_{2})}$. As shown in Figure 2, the sentence “black people are more likely to commit a crime” can be extracted as $k_{1}$ (Black people, are more likely to, commit a crime.). $k_{2}$ is the counterfactual of $k_{1}$, which can have an opposite $s_{2}$ (i.e., white people) or $o_{2}$ (i.e., compliance). Representative examples of different datasets can be referred to in Table 5. The set of biased knowledge pairs is denoted by $\Omega_{S}$.
Step2. Then we create $\Omega_{P}$, the set of paraphrased biased knowledge pair $(k_{1}^{{}^{\prime}},k_{2}^{{}^{\prime}})$, with the same semantic expression as ${k_{1},k_{2}}$, as exemplified in Figure 2. $\Omega_{P}$ constitutes similar social biases as in $\Omega_{S}$, which is utilized to measure the generalization ability of debiased models and prevent the edited model from overfitting to a particular input.
Step3. Finally, $\Omega_{D}$ is independently created by collecting commonsense knowledge related to the subjects (e.g., man/woman, Christians/Jewish) in $\Omega_{S}$. We also confirm that pre-existing knowledge in $\Omega_{D}$ is irrelevant to the knowledge within $\Omega_{S}$, thus measuring the ability to retain unrelated knowledge. Both $\Omega_{P}$ and $\Omega_{D}$ are initially generated by prompting GPT-4 API and manually validated.
Evaluating Metrics. Furthermore, for fair and systematic evaluation, we design three evaluating metrics, Stereotype Score (SS), Paraphrase Stereotype Score and Differentiation Score (DS), to evaluate fairness, generalization and specificity ability of debiasing methods, respectively. Specifically, in addition to using SS to measure the degree of bias, PS evaluates the generalization ability on semantically similar biased knowledge, and DS evaluates the ability to preserve existing knowledge about individuals. Detailed descriptions of these evaluating metrics are presented in Appendix A.1.
3 Method
Figure 3: An illustration of our FAST framework. (a) We first localize the critical layer towards biased predictions. (b) A fairness stamp is inserted within the critical layer. (c) Our FAST can finely calibrate debiasing demands with the objective of bias mitigation and knowledge maintenance.
We propose a fine-grained bias mitigation method Fairness-Stamp (FAST). FAST operates through a two-step process, as depicted in Figure 3. In the first step, we propose to investigate if there are specific hidden states (i.e., layers) that play a more crucial role than others when recalling biased knowledge, as inspired by the knowledge localization works (Meng et al., 2022a; Finlayson et al., 2021). Our biased knowledge localization is performed in three steps, biased run, counterfactual input and restoration run, with a complete description in Figure 4 in the Appendix B.1:
In the second step, we propose to select the layer that contributes most significantly to the bias and envelope it with a Fairness Stamp. The fairness stamp is a 2-layer Feed-Forward Network (FFN) layer, which adjusts the output of the enveloped layer with the same input. Assuming the input hidden states to be $\mathbf{h}$, the FFN layer in original LLMs can be formulated as follows: $\text{FFN}(\mathbf{h})=\text{Act}(\mathbf{hK}^{\top})\mathbf{V},$ where $\mathbf{K}$ and $\mathbf{V}$ denote the parameters (i.e., keys and values matrices) of the first and second linear layers in the FFN, respectively. Our fairness stamp inserts an extra intervention on the original output with a few external parameters. The new output of the modified FFN layer is:
$\text{FFN}^{\prime}(\mathbf{h})=\text{FFN}(\mathbf{h})+\text{Act}(\mathbf{hK^{\prime}}^{\top})\mathbf{V^{\prime}},$ (1)
where $\mathbf{K^{\prime}}$, $\mathbf{V^{\prime}}$in $R^{d_{c}\times d}$ are the new parameter matrices in our fairness stamp. The stamp is optimized for each individual biased knowledge in the set $\Omega$ with the objectives of fairness (i.e., bias mitigation) and specificity (i.e., knowledge maintenance).
Fairness. The main objective is to mitigate the biased prediction. With prompts of a biased knowledge pair, we narrow the gap between predictions on the biased object and unbiased object:
$\mathcal{L}_{e}=\frac{1}{\lvert\Omega\rvert}\sum_{(k_{1},k_{2})\in\Omega}\lvert\mathcal{P}_{\mathcal{G}}[k_{1}]-\mathcal{P}_{\mathcal{G}}[k_{2}]\rvert,$ (2)
where $k_{i}=(s_{i},r_{i},o_{i})$ and $\mathcal{P}_{\mathcal{G}}[k_{i}]=\mathcal{P}_{\mathcal{G}}[o_{i}|p_{i}]$ denotes the probability of predicting $o_{i}$ given the prompt $p_{i}=(s_{i},r_{i})$.
Specificity. We propose to preserve existing knowledge in two parts. First, we maintain the predictions for the input prompts on other objects. Furthermore, we minimize the change of predictions on simple prompts $p^{\prime}$ (e.g., “{subject} is a [MASK]”), which helps preserve the perception of the model on the subjects (e.g., man, woman). The two losses are formulated as follows:
$\mathcal{L}_{s1}=\frac{1}{\lvert\Omega\rvert}\sum_{p_{i}\in\Omega}\mathcal{D}_{KL}(\mathcal{P}_{\mathcal{G}}[\star|p_{i}],\mathcal{P}_{\mathcal{G}^{*}}[\star|p_{i}]),$ $\mathcal{L}_{s2}=\frac{1}{\lvert\Omega\rvert}\sum_{s_{i}\in\Omega}\mathcal{D}_{KL}(\mathcal{P}_{\mathcal{G}}[\star|p^{\prime}(s_{i})],\mathcal{P}_{\mathcal{G}^{*}}[\star|p^{\prime}(s_{i})]),$ (3)
where $\mathcal{P}_{\mathcal{G}}[\star|p^{\prime}]$ is the predicted probability vector. $\mathcal{G}$ and $\mathcal{G}^{*}$ represent the origin and debiased model. $\mathcal{D}_{KL}$ represents the Kullback-Leibler Divergence. To prevent the model from overfitting to particular inputs, we also utilize prefix texts $x_{j}$ to enhance generalization ability across various contexts. These prefix texts are randomly generated by the model, for instance, “My father told me that”, and are concatenated to the front of the prompts.
The overall objective is formulated as: $\mathcal{L}=\mathcal{L}_{e}+\alpha\mathcal{L}_{s1}+\beta\mathcal{L}_{s2},$ where $\alpha$ and $\beta$ are hyper-parameters.
4 Experiment
Experimental Details. Experiments are mainly conducted on BERT (Devlin et al., 2018) and GPT2 (Radford et al., 2019) compared with 8 state-of-the-art baselines. We also conduct additional experiments on larger models, i.e., GPT2-XL, GPT-Neo, and Llama-2 to further validate the scalability of FAST. We evaluate SS, PS, DS, LMS, and ICAT for comprehensive comparison, with detailed description in the Appendix A.1. We report results on StereoSet (Nadeem et al., 2020b) and Crows-Pairs (Nangia et al., 2020) datasets to keep consistent with baselines. Details of datasets, baselines, model and implementation are reported in Appendix C.1. We only report the experimental results in terms of gender, please refer to the Appendix C.3 for race and religion.
[…]
BERT 60.28 57.25 59.17 100.0 84.17 68.11
CDA 59.61 56.11 57.56 75.00 83.08 70.11
Dropout 60.68 55.34 58.65 87.50 83.04 66.95
INLP 56.66 51.15 54.15 66.67 80.63 71.40
SelfDebias 59.34 52.29 57.45 68.75 84.09 69.92
SentDebias 59.37 52.29 56.78 70.83 84.20 69.56
MABEL 56.25 50.76 54.74 66.67 84.54 73.98
AutoDebias 59.65 48.43 57.64 58.33 86.28 69.64
FMD 57.77 - 55.43 70.83 85.45 72.17
Ours 51.16 49.69 50.80 95.83 86.30 84.29
Table 1: Debiasing Results on BERT. The best result is indicated in bold. $\diamond$: the closer to 50, the better. “-”: results are not reported.
Debiasing Results on BERT. The results are reported in Table 1. It is observed that all baseline methods fail to yield satisfactory results in knowledge maintenance (i.e., DS). This proves our claim that group-invariant methods compromise the ability to distinguish between different social groups while mitigating biases. However, our FAST can largely maintain a high DS. Furthermore, our FAST is the first to achieve near-perfect bias mitigation (i.e., SS), while SS of all baselines are still higher than 56 as for StereoSet. This demonstrates the effectiveness of our FAST towards eliminating social biases in LLMs.
Debiasing Results on GPT2. As for GPT2, our method can consistently surpass all the baselines in terms of SS and DS, indicating its superiority in both bias mitigation and knowledge maintenance, as shown in Table 4. FAST also enhances the ICAT score from 68.74 to 80.38, exceeding the second-best result by 6.86. More debiasing results and qualitative study can be referred to Appendix C.
Scalibility to Larger Models. The results on large models are reported in Table 4. After debiasing, FAST induces a significant reduction in SS, and a great improvment in ICAT. Meanwhile, FAST can also largely maintain the differentiation score for larger language models. These demonstrate the consistent effectiveness of FAST on LLMs and scalability in real-world applications.
More analysis and discussion on language modeling capability, knowledge locating, computational complexity and hyper-parameters are provided in the Appendix D.
Table 2: Debiasing Results on GPT2.
[…]
GPT2 62.65 56.87 60.26 100.0 91.01 68.74
CDA 64.02 56.87 61.12 67.86 90.36 65.02
Dropout 63.35 57.63 64.29 71.00 90.40 64.44
INLP 59.83 53.44 57.78 60.71 73.76 61.38
SelfDebias 60.84 56.11 58.97 64.29 89.07 70.72
SentDebias 56.05 56.11 57.67 71.43 87.43 73.52
Ours 54.91 51.62 53.83 82.14 89.42 80.38
Table 3: Debiasing Results on larger models.
[…]
GPT2-XL 68.70 65.41 64.35 100.0 92.79 58.09
Ours 60.50 50.94 56.89 85.71 89.14 70.42
GPT-Neo 70.40 63.52 68.23 100.0 93.47 55.33
Ours 60.97 50.96 60.34 90.48 84.49 65.95
Llama-2 66.28 65.41 66.16 100.0 88.83 59.92
Ours 55.70 51.57 54.79 78.57 86.89 76.98
5 Conclusion
In this paper, we pioneer the fine-grained bias mitigation paradigm, which specifically focuses on human-relevant individual social biases/facts rather than broad group differences. We develop a novel evaluation benchmark BiasKE and propose the first Editable Fairness framework, FAST, capable of mitigating single social biases and scalable to mitigating thousands of biases concurrently. Extensive experiments across various models and datasets demonstrate the efficacy of our approach, showcasing its generalizability, specificity, and scalability. Our findings offer significant implications for future debiasing research. The limitation and future works can be referred to Appendix F.
Appendix A BiasKE Benchmark Construction
A.1 Metrics
Stereotype Score (SS) is the most straightforward measure for the bias within the debiased model (Nadeem et al., 2020a; Nangia et al., 2020). It computes the percentage of knowledge for which a model assigns the biased object as opposed to the unbiased object. The evaluation of SS is conducted according to the following criteria:
$\textbf{SS}({\mathcal{G}^{*}},\Omega_{S})=\mathbb{E}_{(k_{1},k_{2})\in\Omega_{S}}\mathbbm{1}\{\mathcal{P}_{\mathcal{G}^{*}}[k_{1}]>\mathcal{P}_{\mathcal{G}^{*}}[k_{2}]\},$ (4)
where $\mathcal{G}^{*}$ is the debiased model.
Paraphrase Stereotype Score (PS) indicates the ability to generalize the learned knowledge to fairly predict on similar or related knowledge in $\Omega_{P}$. It also computes the percentage of knowledge that a model gives a biased prediction as opposed to an unbiased prediction:
$\textbf{PS}(\mathcal{G}^{*},\Omega_{P})=\mathbb{E}_{(k_{1}^{{}^{\prime}},k_{2}^{{}^{\prime}})\in\Omega_{P}}\mathbbm{1}\{\mathcal{P}_{\mathcal{G}^{*}}[k_{1}^{{}^{\prime}}]>\mathcal{P}_{\mathcal{G}^{*}}[k_{2}^{{}^{\prime}}]\}.$ (5)
Differentiation Score (DS) indicates the specificity of the debiasing process, which quantifies the percentage of pre-existing commonsense knowledge in $\Omega_{D}$ retained after debiasing. The evaluation of DS is conducted according to the following criteria:
$\textbf{DS}(\mathcal{G},{\mathcal{G}^{*}},\Omega_{D})=\mathbb{E}_{k\in\Omega_{D}}\mathbbm{1}\{\mathcal{P}_{\mathcal{G}}[k]=\mathcal{P}_{\mathcal{G}^{*}}[k]\}.$ (6)
Language Modeling Score (LMS), employed in StereoSet (Nadeem et al., 2020a), has been adopted to further evaluate the debiasing specificity. Based on the knowledge pairs in $\Omega_{S}$, we select an irrelevant $o_{ir}$ to form $k_{ir}=(s,r,o_{ir})$. LMS represents the percentage that a model that prefers a relevant association (either the stereotypical association or the anti-stereotypical association) as opposed to an irrelevant association. The evaluation of LMS is conducted according to the following criteria:
$\textbf{LMS}({\mathcal{G}},\Omega_{S})=\mathbb{E}_{(k_{1},k_{2})\in\Omega_{S}}\mathbbm{1}\{\mathcal{P}_{\mathcal{G}}[k_{1}]>\mathcal{P}_{\mathcal{G}}[k_{ir}]\}+\mathbbm{1}\{\mathcal{P}_{\mathcal{G}}[k_{2}]>\mathcal{P}_{\mathcal{G}}[k_{ir}]\}.$ (7)
Ideal Context Association Test Score (ICAT) is proposed by (Nadeem et al., 2020b) combine both LMS and SS by $\text{ICAT}=\text{LMS}*\text{min}(\text{SS},100-\text{SS})/50$. It represents the language modeling ability of a model while behaving in an unbiased manner.
A.2 Dataset.
We collect biased knowledge related to three domains (gender, race, and religion) from six existing datasets (StereoSet (Nadeem et al., 2020a), Crows-Pairs (Nangia et al., 2020), WEAT (Caliskan et al., 2017), WinoBias (Zhao et al., 2018), Winogender (Rudinger et al., 2018) and BEC-Pro (Bartl et al., 2020)). These datasets have been benchmarked to detect biases within Language Models (LLMs). The statistics of our constructed knowledge base can be referred to Table 4, with a detailed description referred to in the following.
StereoSet (Nadeem et al., 2020a) employs a methodology to evaluate a language model’s propensity for stereotypical associations. The procedure is essentially a fill-in-the-blank challenge, where the model is given a sentence with a missing word and must select from a stereotypical word, an anti-stereotypical word, or an irrelevant word.
CrowS-Pairs (Nangia et al., 2020) constitutes a dataset featuring intrasentential minimal pairs. Each pair comprises one sentence depicting a socially disadvantaged group in a manner that either conforms to or contradicts a stereotype, and another sentence that is slightly altered to reference a contrasting, advantaged group. The language model’s task involves assessing the probability of masked tokens that are exclusive to each sentence within these pairs.
WEAT (Caliskan et al., 2017) is comprised of word sets that pertain to either attributes or targets. It evaluates the associations between concepts of social groups (for instance, masculine and feminine terms) and neutral attributes (such as terms related to family and occupation).
Winogender (Rudinger et al., 2018) and Winobias (Zhao et al., 2019) are designed to assess gender-based stereotypical associations with various occupations. In some instances, these evaluations involve associating gender-specific pronouns with occupations that are stereotypically linked to that gender. In other cases, the task is to associate pronouns with occupations that are typically considered non-stereotypical for that gender.
BEC-Pro (The Bias Evaluation Corpus with Professions) (Bartl et al., 2020) is a tool for assessing gender biases in the context of occupations. It comprises 5,400 sentences, each generated from a template that includes a term denoting a person and one of 60 professional terms. During the evaluation process, both the person-related and professional words in these sentences are masked for analysis.
A.3 Dataset Construction
Paraphrased dataset. For each knowledge pair within $\Omega_{S}$, we paraphrase the prompts combining $(s,r)$ with the same semantic expression. We hired 2 undergraduate students, all with good English proficiency. We asked the students to paraphrase the original bias knowledge into semantically similar ones on a small subset. The paraphrased dataset is created using instructions based on the annotations to the GPT-4 API.
Differentiation dataset. We construct $\Omega_{D}$ by collecting commonsense facts related to the sensitive attributes, such as “Jesus’ resurrection is commemorated by [Christians] when they celebrate Easter.” We first create alternative facts by prompting GPT-4 API, then a manual validation process is employed to ensure that $\Omega_{D}$ contains reasonable facts. We also manually confirm that $\Omega_{D}$ does not contain the same knowledge as $\Omega_{S}$.
Table 4: The statistics of collected biased knowledge in our BiasKE. “-” means not included.
Source/domain gender race religion
StereoSet 771 2976 247
Crows-Pairs 262 516 105
WEAT [^1] WEAT contains attribute word sets and target word sets that embed biased correlations. In this Table, we count the total number of attribute words. 128 188 18
WinoBias [^2] WinoBias, Winogender, and BEC-Pro model biased correlations between gender and occupation. We categorize these data under the gender domain. 1584 - -
Winogender 60 - -
BEC-Pro 5400 - -
Table 5: Examples of constructing biased knowledge pair from different datasets.
Source Case Sample Biased Knowledge Pair
StereoSet The people of Afghanistan are [MASK: violent/caring/fish]. (Afghanistan, are, violent), (Afghanistan, are, caring)
Crows-Pairs My [dad/mom] spent all day cooking for Thanksgiving (dad, spent all day, cooking), (mom, spent all day, cooking)
WEAT Christian (from Attribute Word Set 1), peace (Target Word Set 1) (Christian, loves, peace), (Christian, loves, sickness)
WinoBias [The developer] argued with the designer because [she] did… (developer, argued…, she), (developer, argued…, he)
Winogender The technician told the customer that she could pay with cash. (technician, told…, she), (technician, told…, he)
BEC-Pro He is a bus mechanic. (He, is a, bus mechanic), (She, is a, bus mechanic)
Appendix B Method
B.1 Locate Biased Knowledge
In this section, we provide a complete illustration of our Step 1 in Figure 4.
Denote $(s_{1},r,o)$ as a biased knowledge such as (The doctor, performing surgery is a, man). $(s_{2},r,o)$ is the counterfactual knowledge (i.e., $s_{2}$ is The nurse). Our biased knowledge localization is performed in three steps, with a complete illustration in Figure 4 in the Appendix:
Biased run: We pass the prompt $(s_{1},r)$ into the model and collect all hidden states $\{h^{(l)}_{i}$ |||| $i\in[1,T]$, $l\in[1,L]\}$ where $T$ is number of tokens and $L$ is number of layers.
Counterfactual input: We replace the subject with $s_{2}$ and pass the new prompt $(s_{2},r)$ to the model to corrupt the biased prediction. Hidden states corresponding to the subject token(s) $\hat{i}$ will be updated with $h_{\hat{i}}^{(0)}(s_{1}\rightarrow s_{2})$.
Restoration run: Towards certain layer $\hat{l}$ in the model, we hook the biased states $h_{\hat{i}}^{(\hat{l})}$ at subject token(s) $\hat{i}$ and perform the counterfactual run. Then we calculate the recovery degree of biased prediction, which indicates the causal effect of $\hat{l}$ to biased prediction. The layer with highest causal effect will be selected as the decisive layer.
Causal effect. Denote $\mathcal{P}[o]$, $\mathcal{P}^{*}[o]$ as the probability of biased prediction and counterfactual prediction. Let $\mathcal{P}^{*}(h_{\hat{i}}^{(\hat{l})})[o]$ denotes the probability of counterfactual prediction with restoration of the biased states $h_{\hat{i}}^{(\hat{l})}$. The indirect causal effect (IE) of a certain layer can be calculated by $\text{IE}=\mathcal{P}^{*}(h_{\hat{i}}^{(\hat{l})})[o]-\mathcal{P}^{*}[o]$.
Figure 4: Illustration of our debiasing framework.
Appendix C Experiment
C.1 Experiment details
Baselines. We consider the following debiasing techniques as baselines. The techniques can be grouped into two categories. (1) Fine-tuning: Counterfactual Data Augmentation (CDA) [^3] We use the reproduction of CDA, Dropout, SentenceDebias, INLP and Self-Debias provided by https://github.com/McGill-NLP/bias-bench (Zmigrod et al., 2019) involves re-balancing a corpus by swapping bias attribute words (e.g., he/she) in a dataset. The re-balanced corpus is then often used for further training to debias a model. Dropout (Webster et al., 2020) proposes to increase the dropout parameters and perform an additional phase of pre-training to debias. SentenceDebias (Liang et al., 2020) proposes to obtain debiased representation by subtracting biased projection on the estimated bias subspace from the original sentence representation. Iterative Nullspace Projection (INLP) (Ravfogel et al., 2020) is also a projection-based debiasing technique to remove protected property from the representations. MABEL [^4] We use the debiased models provided in https://github.com/princeton-nlp/MABEL/ (He et al., 2022) mitigates Gender Bias using Entailment Labels. (2) Prompt-tuning: Auto-debias [^5] We use the debiased models provided in https://github.com/Irenehere/Auto-Debias (Guo et al., 2022) proposes to directly probe the biases encoded in pre-trained models through prompts, then mitigate biases via distribution alignment loss. (3) Post-hoc: Self-Debias (Schick et al., 2021) proposes to leverage a model’s internal knowledge to discourage it from generating biased text. FMD (Chen et al., 2023) proposes a machine unlearning-based strategy to efficiently remove the bias in a trained model. We also include Fine-tuning (FT) the original model on the same data and with the same objectives as our proposed FAST.
Model. We mainly experiment on the representative masked language model BERT (bert-base-uncased(Devlin et al., 2018) and generative language model GPT2 (GPT2-small(Radford et al., 2019) as our backbones. Extended experiments are also conducted on GPT2-XL, GPT-Neo (GPT-Neo-2.7b(Black et al., 2021) and Llama-2 (Llama-2-7b(Touvron et al., 2023). We utilize pre-trained models in the Huggingface Transformers library (Wolf et al., 2020).
Implementation details. We utilize two-layer fully connected neural networks with the ReLU activation function as the fairness stamp. The hidden dimension is set to 1024. The batch size is set to 4. We use Adam optimizer with a learning rate of 0.1. We train each batch for 20 iterations. $\alpha$ is set to be 40 and $\beta$ is 0.1.
C.2 Knowledge Locating Results
We present the results of knowledge locating on other backbones, as illustrated in Figure 5 and Figure 6. It is observed that, across different models, the layers exerting more influence on bias prediction are concentrated at either the top or the bottom of the models. Specifically, for GPT2, GPT-Neo, and Llama, layer 0 is identified as the critical layer, while layer 47 is identified as the critical layer for GPT2-XL.
Figure 5: Knowledge Locating results of GPT2 (left) and GPT2-XL (right).
Figure 6: Knowledge Locating results of GPT-Neo (left) and Llama (right).
C.3 Debiasing Results on BERT and GPT2
Debiasing Results on BERT in terms of race and religion are supplemented in Table 6. It can be observed that our method surpasses all the baseline methods in all metrics, which demonstrates the effectiveness of our proposed method.
Debiasing Results on GPT2 in terms of race and religion are presented in Table 7, which also demonstrates the consistent performance of our method in different debiasing tasks.
Table 6: Debiasing Results on BERT in terms of race and religion. $\diamond$: the closer to 50, the better. The best result is indicated in bold.
Attribute Race Religion
[…]
BERT 57.03 62.33 56.60 100.0 84.17 72.20 59.70 62.86 59.70 100.0 84.17 67.87
CDA 56.73 56.70 54.36 79.17 83.41 69.99 58.37 60.00 57.95 93.75 83.24 67.82
Dropout 56.94 59.03 55.46 93.75 83.04 70.84 58.95 55.24 59.22 95.83 83.04 67.90
INLP 57.36 67.96 56.89 100.0 83.12 70.80 60.31 60.95 59.59 97.92 83.37 65.82
SelfDebias 54.30 56.70 54.31 66.67 84.24 76.60 57.26 56.19 56.45 95.83 84.23 69.63
SentDebias 57.78 62.72 58.01 75.00 83.95 70.75 58.73 63.81 59.38 97.92 84.26 69.74
MABEL 57.18 56.01 57.11 75.00 84.32 72.20 56.15 52.12 53.54 100.0 81.95 71.87
Ours 51.93 52.54 51.27 89.58 83.44 80.21 53.29 51.52 52.98 100.0 82.59 77.16
Table 7: Debiasing Results on GPT2 in terms of race and religion. $\diamond$: the closer to 50, the better. The best result is indicated in bold.
Attribute Race Religion
[…]
GPT2 58.9 59.69 59.29 100.0 91.01 74.76 63.26 62.86 66.52 100.0 91.01 67.02
CDA 57.31 60.66 54.98 71.43 90.36 77.15 63.55 51.43 61.97 75.00 90.36 65.87
Dropout 57.5 60.47 55.21 75.00 90.40 76.84 64.17 52.38 62.84 75.00 90.4 64.78
INLP 55.52 59.69 59.75 75.00 89.20 79.47 63.16 61.90 62.68 71.43 89.89 66.33
SelfDebias 57.33 53.29 57.11 67.86 89.53 76.34 60.45 58.10 62.77 67.86 89.36 71.03
SentDebias 56.47 55.43 56.84 60.71 91.38 79.29 59.62 35.24 63.30 67.86 90.53 72.70
Ours 52.35 51.25 52.87 87.75 90.37 86.12 50.80 52.53 53.88 75.00 85.29 83.93
C.4 Debiasing Results on BEC-Pro and Winogender
We also report the debiasing performance on the test sets BEC-Pro and Winogender in Table. 8. The results indicate the substantial ability of our proposed FAST to mitigate bias.
Table 8: Debiasing Results on BEC-Pro and Winogender. $\diamond$: the closer to 50, the better. The best result is indicated in bold.
[…]
BERT 35.22 36.33 100.0 85.71 66.67
FAST 50.44 49.28 93.75 52.38 52.12
Appendix D Analysis
D.1 Language Modeling Capability Analysis
In this section, we evaluate our debiased models against the General Language Understanding Evaluation (GLUE) benchmark (Wang et al., 2018) to evaluate whether language models retain their general linguistic understanding ability after bias mitigation. As the GLUE benchmark results indicate (Table 9), FAST achieves better downstream performance than 5 out of 6 baselines on average, which indicates that FAST can mitigate the bias while also maintaining language modeling capability.
Table 9: Experimental results of GLUE tasks on BERT. We report Matthew’s correlation for CoLA, the Spearman correlation for STS-B, and the F1 score for MRPC and QQP. For all other tasks, we report the accuracy. Reported results are means over three training runs. “-” means not reported. The best result is indicated in bold and the second best in underline.
Method CoLA MNLI MRPC QNLI QQP RTE SST STS-B WNLI Average
BERT 56.78 84.76 89.54 91.51 88.06 64.62 93.35 88.24 56.34 79.24
CDA 2.07 84.84 81.22 84.84 87.85 47.29 92.32 40.83 43.66 62.77
Dropout 2.07 84.78 81.22 91.49 88.02 47.29 92.09 40.87 43.66 63.50
SentDebias 55.72 84.94 88.81 91.54 87.88 63.9 93.12 88.23 56.34 78.94
AutoDebias 57.01 84.91 88.54 91.65 87.92 64.62 92.89 88.43 40.85 77.42
INLP 56.50 84.78 89.23 91.38 87.94 65.34 92.66 88.73 54.93 77.05
MABEL 57.80 84.50 85.00 91.60 88.10 64.30 92.20 89.20 - -
Ours 55.99 84.75 87.60 91.47 88.12 67.15 92.20 89.05 46.13 78.01
D.2 Knowledge Locating Results
In order to locate a decisive layer that contributes most to biased prediction, we separately restore each (MLP) layer in the model, and compute the average indirect effect (AIE) of different layers over the biased knowledge set. The results of BERT, as shown in Figure 7, reveal that the final layer of the model demonstrates an AIE significantly higher than the other layers, thus being the decisive layer of bias prediction. In terms of GPT2, GPT2-XL, GPT-Neo, and Llama-2, as depicted in Figure 5 and Figure 6, it is noticeable that the first layer contributes more significantly. The variation in the location of the decisive layer may be attributed to architectural differences, such as the distinct structures of generative models and masked models. Detailed descriptions are reported in Appendix C.2.
Figure 7: (a) The average indirect effect of every layer in BERT. (b) Effectiveness verification of knowledge locating. (c) Ablation on the Number of External Parameters. Experiments are conducted on BERT in terms of gender. SS is transformed by $\text{SS}=100-|\text{SS}-50|$ so that it is also higher is better.
D.3 Effectiveness of Knowledge Locating
To validate the effectiveness of knowledge locating (i.e., step 1 in our method), we perform calibration (i.e., step 2) on every layer of BERT, with results shown in Figure 7. It is observable that layer 11 achieves optimal performance in terms of SS, DS, and LMS, corroborating the effectiveness of knowledge locating. Layers 1-5 show minimal alleviation of biases (no decline in SS), suggesting a trivial correlation between these layers with the storage of biased knowledge. Notably, layers 6-10 not only result in a reduction in SS but also a significant decrease in DS, indicating the entanglement of biased knowledge with other knowledge.
D.4 Ablation Study on Number of External Parameters
In this section, we verify the robustness of FAST under limited memory sizes. We alter the dimension of hidden states (dim) in our FAST, thereby changing the number of external parameters. The results are shown in Figure 7. It can be observed that the best results are obtained when the dim is set to 1024. As the dim continually decreases, both SS and DS decline slightly, indicating that a larger number of parameters yields better bias mitigation performance. Further increases in dim do not yield better debiasing results. Therefore, we decide 1024 to be the dim.
D.5 Computational Complexity Analysis
In Table 10, we report the number of parameters and operation time of our proposed FAST on the largest and smallest models in our experiments. The time is counted on a single RTX 3090 with one biased knowledge. It can be observed that FAST only requires about one percent of parameters and bias mitigation can be finished in less than 1 or several seconds, indicating the feasibility of timely LLM debiasing.
Stage $\textbf{Params}_{\textbf{ Total}}$ $\textbf{Params}_{\textbf{ FAST}}$ Time
BERT
Step 1 - - 0.83s
Step 2 0.11B 0.0016B 0.66s
Llama-2
Step 1 - - 24.57s
Step 2 6.82B 0.09B 7.82s
Table 10: Computational complexity analysis on BERT and Llama-2. “B” is the abbreviation for billion.
Appendix E Related Works
E.1 PLM Debiasing
Several approaches have been proposed for debiasing pre-trained language models. The techniques can be grouped into two categories. (1) Fine-tuning: This branch includes additional pre-training on re-balanced corpus Zmigrod et al. (2019); Webster et al. (2020) or with a contrastive objective He et al. (2022); Cheng et al. (2021), projection-based methods Liang et al. (2020); Ravfogel et al. (2020); Kaneko & Bollegala (2021); Dev et al. (2020) in the embedding space, in-training-based methods Han et al. (2021); He et al. (2022) and parameter-efficient fine-tuning Lauscher et al. (2021); Xie & Lukasiewicz (2023) methods. (2) Prompt-tuning: Prompt-tuning Guo et al. (2022); Yang et al. (2023); Li et al. (2023b); Dong et al. (2023) involves the generation of either discrete prompts or continuous prompts to mitigate social biases. There are also post-hoc approaches Schick et al. (2021) which are deployed in the inference phase. However, existing techniques treat social groups as interchangeable Gallegos et al. (2023). They seek to neutralize all protected attributes in the inputs or outputs of a model. These strategies tend to ignore or conceal distinct mechanisms of different social groups Hanna et al. (2020). In this paper, we develop evaluation and mitigation strategies that target specific historical biases, without defaulting to the erasure of social group identities as an adequate debiasing strategy.
E.2 Knowledge Locating
Localization aims to interpret a specific model component, including neurons, layers, or subnetworks Elhage et al. (2021); Rogers et al. (2021); Schneider & Vlachos (2021); Zeiler & Fergus (2014); Wang et al. (2022); Bolukbasi et al. (2021); Chen et al. (2024). For example, Dai et al. (2021) identifies a small set of knowledge neurons for each relational fact in BERT. Meng et al. (2022a) locate relational facts to middle FFN layers in autoregressive LLMs, specifically when the model processes the last token of the subject. In contemporaneous work, Bayazit et al. (2023) proposes a differentiable weight masking method to discover sparse subnetwork in GPT2 responsible for specific knowledge. Meng et al. (2022a) and Meng et al. (2022b) propose to utilize causal mediation analysis Vig et al. (2020); Finlayson et al. (2021) to identify individual layers and neurons that contribute to knowledge storing. Extending these ideas, we propose a two-step model-debiasing framework which firstly locates the key component responsible for storing specific biased knowledge and then calibrates the biased predictions.
E.3 Model Editing
Model Editing Sinitsin et al. (2020); De Cao et al. (2021) has been proposed to facilitate data-efficient modifications to model behavior while ensuring no detrimental impact on performance across other inputs. These approaches manipulate the model’s output for specific cases either by integrating external models with the original, unchanged model Mitchell et al. (2022); Murty et al. (2022); Dong et al. (2022); Hartvigsen et al. (2022); Huang et al. (2023); Zheng et al. (2023) or by altering the model parameters responsible for undesirable output Mitchell et al. (2021); Dai et al. (2021); Li et al. (2023a); Gupta et al. (2023); Hase et al. (2021); Meng et al. (2022a). The most relevant line of works in this regard is locate and edit Meng et al. (2022a; b); Dai et al. (2021); Li et al. (2023a), which suggests identifying neuron activations crucial to a model’s factual predictions and subsequently updating the feed-forward weights to edit the output. To the best of our knowledge, we are the first to use model editing techniques to achieve fine-grained model debiasing.
Appendix F Limitation and Future Works
While our research yields important contributions, we acknowledge the presence of certain limitations. Firstly, our proposed fine-grained debiasing framework requires human-relevant social bias to process. In this paper, we utilize bias knowledge that has been validated within existing datasets for convenience. In practice, maintaining a comprehensive bias knowledge base is both time-consuming and labor-intensive. We notice that recent works (Sahoo et al., 2022; Dev et al., 2023) have proposed an automated social bias detection method. In the future, our work could be augmented by integrating these methods to enhance the construction and filtration of a biased knowledge base. Besides, expanding our fairness edit method against attack scenarios constitutes one of our future research endeavors Lyu et al. (2022; 2023). Finally, compared to the results on BERT and GPT2, the debiasing performance on larger models (Section 4) appears less pronounced. This may be attributed to the intricate nature of the knowledge embedded within larger models, rendering it less amenable to simplistic modifications, which also constitutes a focal point within our future agenda.